# An ultra-broadband photonic-chip-based traveling-wave parametric amplifier

### v1

@@
Nikolai Kuznetsov Institute of Physics, Swiss Federal Institute of Technology Lausanne (EPFL), CH-1015 Lausanne, Switzerland Center of Quantum Science and Engineering (EPFL), CH-1015 Lausanne, Switzerland    Alberto Nardi IBM Research Europe, Zurich, Säumerstrasse 4, CH-8803 Rüschlikon, Switzerland Institute of Physics, Swiss Federal Institute of Technology Lausanne (EPFL), CH-1015 Lausanne, Switzerland    Alisa Davydova Institute of Physics, Swiss Federal Institute of Technology Lausanne (EPFL), CH-1015 Lausanne, Switzerland Center of Quantum Science and Engineering (EPFL), CH-1015 Lausanne, Switzerland    Mikhail Churaev Institute of Physics, Swiss Federal Institute of Technology Lausanne (EPFL), CH-1015 Lausanne, Switzerland Center of Quantum Science and Engineering (EPFL), CH-1015 Lausanne, Switzerland    Johann Riemensberger johann.riemensberger@ntnu.no Institute of Physics, Swiss Federal Institute of Technology Lausanne (EPFL), CH-1015 Lausanne, Switzerland Center of Quantum Science and Engineering (EPFL), CH-1015 Lausanne, Switzerland Department of Electronic Systems, Norwegian University of Science and Technology, 7491 Trondheim, Norway    Paul Seidler pfs@zurich.ibm.com IBM Research Europe, Zurich, Säumerstrasse 4, CH-8803 Rüschlikon, Switzerland    Tobias J. Kippenberg tobias.kippenberg@epfl.ch Institute of Physics, Swiss Federal Institute of Technology Lausanne (EPFL), CH-1015 Lausanne, Switzerland Center of Quantum Science and Engineering (EPFL), CH-1015 Lausanne, Switzerland
Optical amplification, crucial for modern communication, primarily relies on erbium-doped fiber amplifiers (EDFAs) to enhance signals without distortion. However, EDFAs only cover a portion of the low-loss spectrum of optical fibers. Optical parametric amplification, utilizing material nonlinearities, offers expanded capacity, but has faced limitations in gain and bandwidth. Addressing this, we present a high-gain continuous-traveling-wave parametric amplifier, surpassing the erbium gain window. Leveraging low-loss gallium phosphide-on-insulator photonic integrated waveguides, we achieve up to 35 dB of parametric gain, with a fiber-to-fiber net gain exceeding 10 dB across a wide wavelength range of approximately 140 nm centered at 1550 nm. We demonstrate the capability to amplify ultra-weak signals, extending over 6 orders of magnitude of input powers, optical frequency combs, and communication signals, highlighting the practical potential of our photonic-chip-based amplifier.
@@
Introduction. Optical fibers and the broadband amplification of traveling- and continuous-wave (CW) optical signals have provided pivotal advancements in modern science and technology, particularly in the domain of optical communications for the transmission of information over long distances. A crucial breakthrough for intercontinental optical fiber networks was the introduction of erbium-doped fiber amplifiers (EDFAs) [1, 2], which eliminated the need for frequent electronic signal regeneration and played a pivotal role in the expansion of the World Wide Web. Efforts have been made to extend the bandwidth of the optical gain by using Raman-assisted EDFAs [3] and semiconductor optical amplifiers (SOAs) [4]. In contrast to amplifiers based on rare-earth-doped fibers or the more recently developed rare-earth-doped photonic integrated circuits (PICs) [5, 6], optical parametric amplifiers (OPAs) rely on the intrinsic nonlinearities of optical materials to generate gain. Such traveling-wave parametric amplifiers (TWPAs) were originally developed in the microwave domain, where they, e.g., provide quantum-limited amplification for qubit readout [7].
@@
OPAs have unique properties that set them apart from other means of amplification [8, 9]. The shape of the amplification window is entirely determined by the dispersion and is limited only by the transparency of the material. The central frequency is adjustable, offering flexibility in signal processing. Parametric amplification provides the capability of frequency conversion through the creation of a concomitant idler wave carrying the same information as the input signal but at a different frequency [10, 11]. The reversal of the optical phase of the idler relative to that of the signal can be utilized to compensate for dispersion and mitigate nonlinear effects in transmission systems [12, 13, 14]. Kerr nonlinearity offers a nearly instantaneous response, allowing rapid operation of the OPA. In contrast to the gain from erbium, parametric gain can be adjusted by varying the optical drive power while maintaining a low noise figure, a behavior essential for amplification of weak input signals. Moreover, optical TWPAs operate near the quantum noise limit and offer the ability to perform noiseless phase-sensitive amplification and thus have the potential to increase the span length of long-haul fiber optical links [15, 16]. Lastly, OPAs are unidirectional, making them resistant to optical feedback and parasitic lasing, reducing the need for optical isolators.
@@
To date, time-continuous operation of an OPA with substantial gain has only been demonstrated in fiber-based or bulk-crystal systems, for example, using hundreds of meters of highly nonlinear fibers [17, 18, 8, 19, 20] that require mitigation of strong detrimental Brillouin scattering, or using mechanically cut lithium niobate waveguides [21, 22], for which the $\chi^{(2)}$ nonlinearity requires periodic poling to generate quasi-phase matching and the performance is limited by the available wafer size, as the mode area is large and the bending losses are excessive. In contrast, photonic integrated circuits (PICs) offer the possibility of a significantly increased parametric gain coefficient $g=\sqrt{(\gamma P_{\textrm{P}})^{2}-(\kappa/2)^{2}}$ [8, 23], which is determined by the effective nonlinearity $\gamma={n_{2}\omega_{\textrm{P}}}/{cA_{\mathrm{eff}}}$ and the phase-mismatch parameter $\kappa$; here, $n_{2}$ is the Kerr nonlinearity, $A_{\mathrm{eff}}$ is the effective mode area, $\omega_{\textrm{P}}$ is the pump angular frequency, and $P_{\textrm{P}}$ is the pump power. With the proper choice of material, PICs can exhibit a strong Kerr nonlinearity while simultaneously featuring a high linear refractive index that substantially reduces the mode area. The material should also have negligible two-photon absorption and a high optical-damage threshold. Integrated photonic platforms that have been studied so far include silicon [24, 25, 26, 27, 28], chalcogenides [29], Si7N3 [30], AlGaAs [31], and highly doped silica [32]. None of these materials have yet produced net gain in the continuous pumping regime, as required for time-varying signals used in most applications. Recently, triggered by the advances in ultra-low-loss meter-long Si3N4 integrated photonic waveguides [33], it was possible to achieve CW parametric amplification with net gain [34, 35]. However, this approach still requires coping with high pump-power levels, meter-long waveguides, limited gain, and narrow bandwidth. Here, we address these challenges by demonstrating a compact, ultra-broadband, high-gain, PIC-based optical TWPA.
Results. We use thin-film gallium phosphide (GaP) on silicon dioxide to create a TWPA comprising a dispersion engineered waveguide operating at a pump wavelength near 1550 nm (see the Supplementary Material for details of the fabrication). GaP combines a high optical refractive index $(n=3.05)$ with a strong Kerr nonlinearity ($n_{2}=1.1\times 10^{-17}$ m2 W-1) and an indirect bandgap ($E_{\mathrm{g}}=2.24$ eV)  [36] sufficiently large to mitigate two-photon absorption at telecommunication wavelengths. This exceptional confluence of properties has facilitated the generation of low-threshold frequency combs [37] and dissipative Kerr solitons [38]. We estimate the effective nonlinearity of the GaP waveguides in this work to be $\gamma=165$ W-1m-1, which is more than 300 times larger than the value of 0.51 W-1m-1 reported for Si3N4 waveguides [35], facilitating a 35-fold reduction in waveguide length and a 60-fold reduction in device footprint. Advances in fabrication techniques [37, 38] allow us to reduce GaP waveguide optical propagation losses below 1 dB cm-1 and achieve a high bonding yield and low defect density. These developments are pivotal for successful fabrication of cm-long waveguide spirals with a footprint of only 500 $\times$ 500 $\mu$m2 (Figs. 1(C - G) & Supplementary Material).
We operate the TWPA with a single pump to amplify signals via degenerate four-wave-mixing (FWM), exploiting the optical Kerr effect (Fig. 1(a)). The optical energy is redistributed by annihilating two pump photons to amplify signal photons at a separate frequency, while simultaneously generating a phase-conjugated idler at a frequency equidistant from the pump laser but offset in the opposite direction. The near instantaneous action of the optical Kerr effect necessitates phase-matching in the waveguide for efficient amplification [8], i.e., $\kappa=\Delta\beta+2\gamma P_{\textrm{P}}$ is small, where $\Delta\beta$ is the linear propagation mismatch as given by $\Delta\beta\approx\beta_{2}\left(\omega_{\textrm{S}}-\omega_{\textrm{P}}\right)^{2}+\beta_{4}\left(\omega_{\textrm{S}}-\omega_{\textrm{P}}\right)^{4}/12+...$, $\beta_{2}$ and $\beta_{4}$ are the second- and fourth-order derivatives with respect to the angular frequency $\omega$ of the optical propagation constant $\beta$ evaluated at the pump frequency $\omega_{\textrm{P}}$, and $\omega_{\textrm{S}}$ is the signal angular frequency. Only even order dispersion terms contribute to the shape of the amplification spectrum. The largest optical bandwidth is achieved when the dispersion is anomalous ($\beta_{2}<0$) and the linear phase mismatch $\Delta\beta$ is compensated by the nonlinear phase mismatch $2\gamma P_{\textrm{P}}$ originating from self- and cross-phase modulation. The peak gain in a waveguide of length $L$ is defined by $G_{\textrm{S}}=1+\left[\sinh(-\Delta\beta L_{\textrm{eff}}/2)\right]^{2}$, where $L_{\textrm{eff}}=(1-\exp(-\alpha L))/\alpha$ is the effective length and $\alpha$ is the linear propagation loss. Fig. 1(b) shows a comparison of the gain spectrum calculated for an optimized GaP optical TWPA with that of C-band and L-band EDFAs. The optimized GaP waveguide has a cross-section of $789\times 299$ $\text{nm}^{2}$, for which $\beta_{2}=-16$ fs2mm-1 and $\beta_{4}=3547$ fs4mm-1. The region of signal and idler gain extends from around 1300 nm in the optical O-band to 1900 nm, well beyond the longest wavelengths used in silica fiber communications. We find that the waveguides used in this work exhibit a stronger anomalous dispersion than expected from our simulations (see Supplementary Material).
@@
Figure 1: Optical continuous-wave parametric amplification in a integrated GaP photonic waveguide. (a) Principle of broadband optical parametric amplification in an integrated spiral waveguide. Strong pump and small signal waves co-propagate through the waveguide, leading to the amplification of the signal and generation of the third wave, the idler. The frequency of the signal wave can be chosen over a wide range within the amplification bandwidth. (b) Comparison of the typical amplification bandwidth of state-of-the-art EDFAs (green: C-band, yellow: L-band) with the amplification bandwidth of integrated GaP parametric amplifiers (red: achieved in this work, blue: if further optimized). (c) Optical microscope image of the fabricated photonic chip with multiple spiral waveguides and other test structures. (d) Enlarged optical microscope image of a 5-cm-long spiral waveguide. Accounting for the chip width of 0.55 cm, the total waveguide length is 5.55 cm. (e) Scanning electron microscope image of a few waveguide turns. (f) Scanning electron microscope image of an inverse taper facet, similar to the one used in this work, used to couple light into a GaP waveguide. (g) Material stack of the fabricated chip.
We measure the CW amplification spectrum of a 5.55 cm-long GaP spiral waveguide with a height of 299 nm and a design width of 780 nm using the same methods as described in Ref. [35]. Figs. 2(A, B) depict the experimental setup and recording scheme used to measure the parametric gain (see Supplementary Material). Fig. 2(c) shows the power-calibrated amplification spectrum. In the spectral region around the pump wavelength, we observed a broadened peak due to the leakage of the amplified spontaneous emission from the EDFA installed in the optical path of the pump; our tunable filter has a bandwidth of 1 nm and a finite suppression ratio of 30 dB. The frequency conversion and amplification spectra (Figs. 2(D, E)) are determined by comparing the output spectra recorded as the the signal laser is swept at various pump power levels (red to orange) with the output spectrum obtained when the pump is switched off (black). Increasing the input power from 1.00 W to 4.13 W, we observe increasing gain, conversion efficiency, and amplification bandwidth, following established theory [39, 8]. The generated idler spectrum spans from 1550 nm to approximately 1480 nm at the highest pump power. In Fig. 2(c), spurious light generated outside the scan range of the signal laser (outer shaded regions) originates from non-degenerate FWM. Hence, we only calculate the internal idler conversion efficiency (Fig. 2(d)) from 1480 nm to 1540 nm and the internal signal gain from 1560 nm to 1630 nm (Fig. 2(e)). We separately measure the passive transmission loss of the TWPA, i.e., the insertion loss, by bypassing the lensed fibers and the chip; the ratio of the spectrum obtained with the bypass to the spectrum transmitted through the chip is plotted as a black line in Fig. 2(e). To calculate the off-chip net gain, we subtract this 10 – 12 dB of insertion loss, revealing net-gain of up to $G_{\textrm{S}}=25$ dB, with $G_{\textrm{S}}>10$ dB over a bandwidth of 70 nm. Assuming a symmetric gain spectrum on the short-wavelength side of the pump, we estimate the 10 dB off-chip gain region to exceed 140 nm, or 17 THz. This bandwidth is almost ten times larger than reported in prior work on Si3N4 [35], and nearly three times larger than the bandwidth of a single C-band EDFA.
@@
Figure 2: Observation of a broadband continuous wave parametric amplification in a GaP spiral waveguide. (a) Schematic of the experimental setup. EDFA: erbium-doped fiber amplifier, (d)VOA: (digital) variable optical attenuator, PC: polarization controller, PM: power meter, OSA: optical spectrum analyzer. (b) Principle of the "Max Hold" setting of the OSA used in broadband gain measurements. With every new scan, the OSA records and updates only the highest values across the measurement span, while the wavelength of the signal laser is slowly swept to cover the whole amplification bandwidth. (c) Measured amplification spectra for different off-chip pump powers. The resolution bandwidth of the OSA is 2 nm. EDFA and DVOA in the signal path are not used in this experiment. (d, e) Calculated idler conversion efficiency (CE) and signal amplification spectra for fiber-coupled input pump powers up to 4.13 W, respectively. (f) Off-chip signal power at wavelength 1605 nm at the output of the amplifier as a function of signal and pump powers at the input fiber. The dashed line shows the boundary when the input signal power is equal to the output signal power. (g) Measurement of the optical TWPA noise figure as a function of signal input power. The dashed line indicates the 3 dB quantum limit of a phase-insensitive amplifier.
To investigate the maximum signal output power and the saturation power of the amplifier, we perform a power sweep of the signal at a wavelength of 1605 nm (Fig. 2(f)). Here we use a different spiral with a designed waveguide width of 790 nm; this spiral is used also in the remaining experiments. The measurements are carried out with the resolution bandwidth of the OSA set at 0.1 nm. At a pump power of 4.43 W, the saturated off-chip signal output power exceeds 125 mW, corresponding to a power of 220 mW on chip at the waveguide output. Hence, the on-chip power conversion efficiency reaches 9 %, which is limited by the waveguide propagation losses. Here, we also demonstrate the ability to achieve low-noise amplification with high gain over a large dynamic range. The linear, small-signal regime of amplification extends up to a signal input power of almost 0 dBm at the highest pump power level of 4.43W. At 2.01 W pump power, the amplification is linear over the entire range of input powers measured, extending over six orders of magnitude and limited only by the capabilities of our measurement setup; in principle, parametric amplification poses no lower limit on the signal input power. We calculate the amplifier noise figure from the measured signal amplification and the spontaneous parametric fluorescence background (see Supplementary Material) [40]. The off-chip noise figure approaches 6 dB in the small-signal gain regime. Since degradation of the signal-to-noise ratio (SNR) at the input facet is irreversible, we take into account the coupling losses of 2.5 dB at the chip input facet and find the on-chip noise figure of less than 4 dB for a wide range of signal powers below saturation (Fig. 2(g)). The degradation of the noise figure at higher signal input powers is attributed to gain saturation, in other words, the signal gain becomes weaker while the parametric fluorescence is still within the small-signal regime and undergoes a higher amplification.

### v2

@@
thanks: These authors contributed equally to this work.thanks: These authors contributed equally to this work.
@@
Nikolai Kuznetsov Institute of Physics, Swiss Federal Institute of Technology Lausanne (EPFL), CH-1015 Lausanne, Switzerland Center of Quantum Science and Engineering (EPFL), CH-1015 Lausanne, Switzerland    Alberto Nardi Institute of Physics, Swiss Federal Institute of Technology Lausanne (EPFL), CH-1015 Lausanne, Switzerland IBM Research Europe, Zurich, Säumerstrasse 4, CH-8803 Rüschlikon, Switzerland    Johann Riemensberger Institute of Physics, Swiss Federal Institute of Technology Lausanne (EPFL), CH-1015 Lausanne, Switzerland Center of Quantum Science and Engineering (EPFL), CH-1015 Lausanne, Switzerland Present address: Department of Electronic Systems, Norwegian University of Science and Technology, 7491 Trondheim, Norway    Alisa Davydova Institute of Physics, Swiss Federal Institute of Technology Lausanne (EPFL), CH-1015 Lausanne, Switzerland Center of Quantum Science and Engineering (EPFL), CH-1015 Lausanne, Switzerland    Mikhail Churaev Institute of Physics, Swiss Federal Institute of Technology Lausanne (EPFL), CH-1015 Lausanne, Switzerland Center of Quantum Science and Engineering (EPFL), CH-1015 Lausanne, Switzerland    Paul Seidler pfs@zurich.ibm.com IBM Research Europe, Zurich, Säumerstrasse 4, CH-8803 Rüschlikon, Switzerland    Tobias J. Kippenberg tobias.kippenberg@epfl.ch Institute of Physics, Swiss Federal Institute of Technology Lausanne (EPFL), CH-1015 Lausanne, Switzerland Center of Quantum Science and Engineering (EPFL), CH-1015 Lausanne, Switzerland Institute of Electrical and Micro Engineering (IEM), Swiss Federal Institute of Technology Lausanne (EPFL), CH-1015 Lausanne, Switzerland
Optical amplification, crucial for modern communication and data center interconnects, primarily relies on erbium-doped fiber amplifiers (EDFAs) to enhance signals without distortion [1, 2]. While EDFAs were historically decisive for the introduction of dense wavelength-division multiplexing, they only cover a portion of the low-loss spectrum of optical fibers. This has motivated the development of amplifiers that operate beyond the erbium gain window, such as Raman amplifiers. Pioneering work on optical traveling-wave parametric amplifiers (TWPAs) [3, 4] utilizing intrinsic third-order optical nonlinearity has led to demonstrations of increased channel capacity and performance. TWPAs offer high gain, can reach the 3-dB quantum limit for phase-preserving amplifiers, and exhibit unidirectional operation. Yet, despite the use of highly nonlinear fibers [5, 6, 3, 7, 8] or bulk waveguides [9, 10], their power requirements and technical complexity have impeded adoption. In contrast, TWPAs based on integrated photonic circuits offer the advantages of substantially increased mode confinement and optical nonlinearity but have been limited in bandwidth because of the trade-off with maintaining low propagation loss [11, 12]. We overcome this challenge by using low-loss gallium phosphide-on-silicon dioxide [13, 14, 15] photonic integrated circuits and attain up to 35 dB of parametric gain with waveguides only a few centimeters long in a compact footprint of only 0.25 square millimeters. Fiber-to-fiber net gain exceeding 10 dB across an ultra-broad bandwidth of approximately 140 nm (i.e., 17 THz) is achieved, surpassing the gain window of a standard C-band EDFA by more than 3 times. We furthermore demonstrate the capability to handle ultra-weak signals; input powers can range over six orders of magnitude while maintaining a low noise figure. We exploit these performance characteristics to amplify both optical frequency combs and coherent communication signals. This marks the first ultra-broadband, high-gain, continuous-wave amplification in a photonic integrated circuit, opening up new capabilities for next-generation interconnects in data centers, artificial-intelligence accelerators, and high-performance computing, as well as optical communication, metrology, and sensing.
@@
Introduction Optical fibers and the broadband amplification of traveling- and continuous-wave (CW) optical signals have provided pivotal advancements in modern science and technology, particularly in the domain of optical communications for the transmission of information over long distances. A crucial breakthrough for intercontinental optical fiber networks was the introduction of erbium-doped fiber amplifiers (EDFAs) [1, 2] that could simultaneously amplify multiple wavelengths of light and eliminate the need for frequent electronic signal regeneration. EDFAs hence played a decisive role in the expansion of the World Wide Web. Efforts have been made to extend the bandwidth of the optical gain by using Raman-assisted EDFAs [16] and semiconductor optical amplifiers (SOAs) [17]. In contrast to amplifiers based on rare-earth-doped fibers or the more recently developed rare-earth-doped photonic integrated circuits (PICs) [18, 19], optical parametric amplifiers (OPAs) rely on the intrinsic nonlinearities of optical materials to generate gain. Such traveling-wave parametric amplifiers (TWPAs) were originally developed in the microwave domain, where they, e.g., provide quantum-limited amplification for qubit readout [20].
@@
OPAs have unique properties that set them apart from other means of amplification [3, 4]. The shape of the amplification window is entirely determined by the optical dispersion and is limited only by the transparency of the material and its nonlinear absorption threshold. The central frequency is adjustable, offering flexibility in signal processing. Parametric amplification provides the capability of frequency conversion through the creation of a concomitant idler wave carrying the same information as the input signal but at a different frequency [21, 22]. The reversal of the optical phase of the idler relative to that of the signal can be utilized to compensate for dispersion and mitigate nonlinear effects in transmission systems [23, 24, 25]. Kerr nonlinearity provides a nearly instantaneous response, allowing rapid operation of the OPA. In contrast to the gain from erbium, parametric gain can be adjusted in situ by varying the optical drive power while maintaining a low noise figure, a behavior essential for amplification of weak input signals. Moreover, optical TWPAs operate near the quantum noise limit and offer the ability to perform noiseless phase-sensitive amplification and thus have the potential to increase the span length of long-haul fiber-optic links [26, 27]. Lastly, OPAs are unidirectional, making them resistant to optical feedback and parasitic lasing, reducing the need for optical isolators in the optical path after the parametric amplifier, particularly in systems reliant on cascaded amplification chains. The increased insensitivity to reflections from the chip facet suggests significant suppression of parasitic lasing within the OPA system, which can readily occur with high on-chip gain. Although it is currently not feasible to remove the isolator from the EDFAs typically used to achieve high pump powers in OPA systems, they could be replaced by semiconductor laser diodes in the future.
@@
To date, time-continuous operation of an OPA with substantial gain has only been demonstrated in fiber-based or bulk-crystal systems. For example, OPAs can be constructed by employing hundreds of meters of highly nonlinear fibers [5, 6, 3, 7, 8], but they require mitigation of strong detrimental Brillouin scattering. Alternatively, mechanically cut lithium niobate waveguides [9, 10] can be used, for which the $\chi^{(2)}$ nonlinearity requires periodic poling to generate quasi-phase matching and the performance is limited by the available wafer size, as the mode area is large and bending losses are excessive. In contrast, photonic integrated circuits (PICs) and particularly integrated waveguides, which are equivalent to fiber OPAs, offer the possibility of a significantly increased parametric gain coefficient $g=\sqrt{(\gamma P_{\textrm{P}})^{2}-(\kappa/2)^{2}}$ [3, 28], which is determined by the effective nonlinearity $\gamma={n_{2}\omega_{\textrm{P}}}/{cA_{\mathrm{eff}}}$ and the phase-mismatch parameter $\kappa$; here, $n_{2}$ is the Kerr nonlinearity, $A_{\mathrm{eff}}$ is the effective mode area, $\omega_{\textrm{P}}$ is the pump angular frequency, and $P_{\textrm{P}}$ is the pump power. With the proper choice of material, PICs can exhibit a strong Kerr nonlinearity while simultaneously featuring a high linear refractive index that substantially reduces the mode area. The material should also have negligible two-photon absorption and a high optical-damage threshold. Integrated photonic platforms that have been studied so far include silicon [29, 30, 31, 32, 33], chalcogenides [34], Si7N3 [35], AlGaAs [36], and highly doped silica [37]. None of these materials have yet produced net gain in the continuous pumping regime, as required for time-varying signals used in most applications. Recently, triggered by advances in ultra-low-loss Si3N4 integrated photonic waveguides [38], it was possible to achieve CW parametric amplification with net gain [12, 11]. However, this approach still requires coping with high pump-power levels, meter-long waveguides, limited gain, and narrow bandwidth. Here, we address these challenges by demonstrating a compact, ultra-broadband, high-gain, PIC-based optical TWPA.
Optical gain measurements We use thin-film gallium phosphide (GaP) on silicon dioxide to create a TWPA comprising a dispersion engineered waveguide operating at a pump wavelength near 1550 nm (see Supplementary Material for details of the fabrication).
@@
Figure 1: Optical continuous-wave parametric amplification in a integrated GaP photonic waveguide. (a) Principle of broadband optical parametric amplification in an integrated spiral waveguide. Strong pump and small signal waves co-propagate through the waveguide, leading to the amplification of the signal and generation of the third wave, the idler. The frequency of the signal wave can be chosen over a wide range within the amplification bandwidth. (b) Comparison of the typical amplification bandwidth of state-of-the-art EDFAs (green: C-band, yellow: L-band) with the amplification bandwidth of integrated GaP parametric amplifiers (red: achieved in this work, blue: if further optimized). Gain values are provided for reference only, and not meant for direct comparison. (c) Optical microscope image of the fabricated photonic chip with multiple spiral waveguides and other test structures. (d) Enlarged optical microscope image of a 5-cm-long spiral waveguide. Accounting for the chip width of 0.55 cm, the total waveguide length is 5.55 cm. (e) Scanning electron microscope image of a few waveguide turns. (f) Scanning electron microscope image of an inverse taper facet, similar to the one used in this work, used to couple light into a GaP waveguide. (g) Material stack of the fabricated chip.
GaP combines a high optical refractive index $(n=3.05)$ with a strong Kerr nonlinearity ($n_{2}=1.1\times 10^{-17}$ m2 W-1) and an indirect bandgap [39] ($E_{\mathrm{g}}=2.24$ eV), sufficiently large to mitigate two-photon absorption at telecommunication wavelengths. This exceptional confluence of properties has facilitated the generation of low-threshold frequency combs [14] and dissipative Kerr solitons [15]. We estimate the effective nonlinearity of the GaP waveguides in this work to be $\gamma=165$ W-1m-1, which is more than 300 times larger than the value of 0.51 W-1m-1 reported for Si3N4 waveguides [11], facilitating a 35-fold reduction in waveguide length and a 60-fold reduction in device footprint. Advances in fabrication techniques [14, 15] allow us to reduce GaP waveguide optical propagation losses to 0.8 dB cm-1 on average within the S-, C- and L-band (see Supplementary Material), and achieve a high bonding yield and low defect density. These developments are pivotal for successful fabrication of centimeter-long waveguide spirals with a footprint of only 500 $\times$ 500 $\mu$m2 (Figs. 1(c - g) and Supplementary Material).
@@
Figure 2: Observation of a broadband continuous-wave parametric amplification in a GaP spiral waveguide. (a) Schematic of the experimental setup. EDFA: erbium-doped fiber amplifier; (D)VOA: (digital) variable optical attenuator; PC: polarization controller; PM: power meter; OSA: optical spectrum analyzer. (b) Principle of the "Max Hold" setting of the OSA used in broadband gain measurements. With every new scan, the OSA records and updates only the highest values across the measurement span, while the wavelength of the signal laser is slowly swept to cover the whole amplification bandwidth. (c) Measured amplification spectra for various off-chip pump powers. The resolution bandwidth of the OSA is 2 nm. EDFA and DVOA in the signal path are not used in this experiment. (d, e) Calculated idler conversion efficiency (CE) and signal amplification spectra for various input pump powers up to 4.13 W. (f) Off-chip signal power at wavelength 1605 nm at the output of the amplifier as a function of signal and pump powers at the input fiber. The resolution bandwidth of the OSA is 0.1 nm. The dashed line is the boundary at which the input signal power is equal to the output signal power. (g) Measurement of the optical TWPA noise figure as a function of signal input power. The dashed line indicates the 3 dB quantum limit of a phase-insensitive amplifier.
We operate the TWPA with a single pump to amplify signals via degenerate four-wave-mixing (DFWM), exploiting the optical Kerr effect (Fig. 1(a)). The optical energy is redistributed by annihilating two pump photons to amplify signal photons at a separate frequency, while simultaneously generating a phase-conjugated idler at a frequency equidistant from the pump laser but offset in the opposite direction. The near instantaneous action of the optical Kerr effect necessitates phase-matching in the waveguide for efficient amplification [3], i.e., $\kappa=\Delta\beta+2\gamma P_{\textrm{P}}$ is small, where $\Delta\beta$ is the linear propagation mismatch given by $\Delta\beta\approx\beta_{2}\left(\omega_{\textrm{S}}-\omega_{\textrm{P}}\right)^{2}+\beta_{4}\left(\omega_{\textrm{S}}-\omega_{\textrm{P}}\right)^{4}/12+...$, $\beta_{2}$ and $\beta_{4}$ are the second- and fourth-order derivatives with respect to the angular frequency $\omega$ of the optical propagation constant $\beta$ evaluated at the pump frequency $\omega_{\textrm{P}}$, and $\omega_{\textrm{S}}$ is the signal angular frequency. Only even order dispersion terms contribute to the shape of the amplification spectrum. The largest optical bandwidth is achieved when the dispersion is slightly anomalous ($\beta_{2}<0$) and the linear phase mismatch $\Delta\beta$ is compensated by the nonlinear phase mismatch $2\gamma P_{\textrm{P}}$ originating from self- and cross-phase modulation. The peak gain in a waveguide of length $L$ is defined by $G_{\textrm{S}}=1+\left[\sinh(-\Delta\beta L_{\textrm{eff}}/2)\right]^{2}$, where $L_{\textrm{eff}}=(1-\exp(-\alpha L))/\alpha$ is the effective length and $\alpha$ is the linear propagation loss. Fig. 1(b) shows a comparison of the gain spectrum calculated for an optimized GaP optical TWPA with that of C-band and L-band EDFAs. The optimized GaP waveguide has a cross-section of $789\times 299$ $\text{nm}^{2}$, for which $\beta_{2}=-16$ fs2mm-1 and $\beta_{4}=3547$ fs4mm-1. The region of signal and idler gain extends from around 1300 nm in the optical O-band to 1900 nm, well beyond the longest wavelengths used in silica fiber communications. We find that the waveguides used in this work exhibit a stronger anomalous dispersion, $\beta_{2}=-124$ fs2mm-1, than expected from our simulations (see Supplementary Material). We measure the CW amplification spectrum of a 5.55 cm-long GaP spiral waveguide with a height of 299 nm and a design width of 780 nm using the same methods as described in Ref. [11]. Figures. 2(a, b) depict the experimental setup and recording scheme used to measure the parametric gain (see Supplementary Material), and Fig. 2(c) shows the power-calibrated amplification spectrum. In the spectral region around the pump wavelength, we observe a broadened peak due to the leakage of the amplified spontaneous emission from the EDFA installed in the optical path of the pump; our tunable filter has a bandwidth of 1 nm and a finite suppression ratio of 30 dB. The frequency conversion and amplification spectra (Figs. 2(d, e)) are determined by comparing the output spectra recorded as the signal laser (set to 1.5 $\mu$W) is swept at various pump power levels (red to orange) with the output spectrum obtained when the pump is switched off (black). Increasing the input power from 1.00 W to 4.13 W, we observe increasing gain, conversion efficiency, and amplification bandwidth, following established theory [40, 3]. The exponential increase in gain with a linear change in pump power is observed, as expected. The generated idler spectrum spans from 1550 nm to approximately 1480 nm at the highest pump power. In Fig. 2(c), spurious light generated outside the scan range of the signal laser (outer shaded regions) originates from non-degenerate FWM. Hence, we only calculate the internal idler conversion efficiency (Fig. 2(d)) from 1480 nm to 1540 nm and the internal signal gain from 1560 nm to 1630 nm (Fig. 2(e)). We separately measure the passive transmission loss of the TWPA, i.e., the insertion loss, by bypassing the lensed fibers and the chip; the ratio of the spectrum obtained with the bypass to the spectrum transmitted through the chip is plotted as a black line in Fig. 2(e). To calculate the off-chip net gain, we subtract this 10 – 12 dB of insertion loss, revealing net-gain of up to $G_{\textrm{S}}=25$ dB, with $G_{\textrm{S}}>10$ dB over a bandwidth of 70 nm. Assuming a symmetric gain spectrum on the short-wavelength side of the pump, we estimate the 10 dB off-chip gain region to exceed 140 nm, or 17 THz. This bandwidth is almost ten times larger than reported in prior work on Si3N4 [11], and nearly three times larger than the bandwidth of a single C-band EDFA. To investigate the maximum signal output power and the saturation power of the amplifier, we perform a power sweep of the signal at a wavelength of 1605 nm (Fig. 2(f)). Here we use a different spiral with a designed waveguide width of 790 nm; this spiral is used also in the remaining experiments. The measurements are carried out with the resolution bandwidth of the OSA set at 0.1 nm. At a pump power of 4.43 W, the saturated off-chip signal output power exceeds 125 mW, corresponding to a power of 220 mW on chip at the waveguide output. Hence, the on-chip power conversion efficiency reaches 9 %, which is limited by the waveguide propagation losses. Here, we also demonstrate the ability to achieve low-noise amplification with high gain over a large dynamic range. The linear, small-signal regime of amplification extends up to a signal input power of almost 0 dBm at the highest pump power level of 4.43 W.